Show that the Moment generating function for a Bernoulli (p) random variable X is given by



MX(t) = pet + q where q = 1 − p.




1
Expert's answer
2021-11-17T03:46:34-0500

Solution:

From the definition of the Bernoulli distribution, X has probability mass function:

"\operatorname{Pr}(X=n)= \begin{cases}q & : n=0 \\ p & : n=1 \\ 0 & : n \notin\{0,1\}\end{cases}"

From the definition of a moment generating function:

"M_{X}(t)=\mathrm{E}\left(e^{t X}\right)=\sum_{n=0}^{1} \operatorname{Pr}(X=n) e^{t n}"

So:

"\begin{aligned}

M_{X}(t) &=\operatorname{Pr}(X=0) e^{0}+\operatorname{Pr}(X=1) e^{t} \\

&=q+p e^{t}

\end{aligned}"

where "q=1-p"

Hence, proved.